OpenStudy (anonymous):

evaluate:∫4dt/(5t+2) i need also the solution...tnx,.

OpenStudy (amistre64):

you need 5 instead of 4 up top; so modify it as needed

OpenStudy (anonymous):

amistre will help you find the antiderivative

OpenStudy (amistre64):

\[\int\frac{4}{(5t+2)}dt\] \[\int\frac{5}{5}*\frac{4}{(5t+2)}dt\] \[\frac{4}{5}\int\frac{5}{(5t+2)}dt\implies\frac{4}{5}ln(5t+2)\]

OpenStudy (anonymous):

use u = 5t +2 you will get \[\ln(5t+2) \] but you need to adjust the constant

OpenStudy (zarkon):

Can I add to your solution... \[\frac{4}{5}\ln|5t+2|+c\]

OpenStudy (amistre64):

..... lol + C

OpenStudy (zarkon):

and \[|\cdot |\] ;)

OpenStudy (anonymous):

thank you very much to all of u for the suggestions and answer...god bless...